Given that f(x) = 5x - 10 and g(x) = x + 3, solve for f(g(x)) when x = -1.
Answer by jim_thompson5910(35256)

f(x) = 5x - 10

f(g(x)) = 5( g(x) ) - 10

f(g(x)) = 5( x+3 ) - 10

f(g(x)) = 5x+15 - 10

f(g(x)) = 5x+5

f(g(-1)) = 5(-1)+5

f(g(-1)) = -5+5

f(g(-1)) = 0